Write down the parametric equations of the following ellipses. Find their eccentricities, foci and directrices. (a) 4x2 + 9y2 = 4 (b) x2/9 + y2/16 = 1 (c) x2/7 + y2/14 = 1


1
Expert's answer
2021-08-13T12:48:18-0400

(a)

"4x^2+9y^2=4"

"\dfrac{x^2}{1}+\dfrac{y^2}{4/9}=1"

"x=\cos t, y=\dfrac{2}{3}\sin t, 0\leq t\leq 2\pi"

"\vec r=\langle\cos t, \dfrac{2}{3}\sin t\rangle, 0\leq t\leq 2\pi"

"a=1, b=\dfrac{2}{3}, c=\sqrt{a^2-b^2}=\sqrt{(1)^2-(\dfrac{2}{3})^2}=\dfrac{\sqrt{5}}{3}"


"e=\dfrac{c}{a}=\dfrac{\sqrt{5}}{3}"

"Foci: (-\dfrac{\sqrt{5}}{3}, 0), (\dfrac{\sqrt{5}}{3}, 0)"

"Directrices: x=\pm\dfrac{a^2}{c}"

"first\ directrix: x=-\dfrac{3\sqrt{5}}{5},"

"second\ directrix:x=\dfrac{3\sqrt{5}}{5}"

(b)

"\dfrac{x^2}{9}+\dfrac{y^2}{16}=1"

"x=3\cos t, y=4\sin t, 0\leq t\leq 2\pi"

"\vec r=\langle3\cos t, 4\sin t\rangle, 0\leq t\leq 2\pi"

"a=4, b=3, c=\sqrt{a^2-b^2}=\sqrt{(4)^2-(3)^2}=\sqrt{7}"


"e=\dfrac{c}{a}=\dfrac{\sqrt{7}}{4}"

"Foci: (-\sqrt{7}, 0), (\sqrt{7}, 0)"

"Directrices: y=\pm\dfrac{a^2}{c}"

"first\ directrix: y=-\dfrac{16\sqrt{7}}{7},"

"second\ directrix:y=\dfrac{16\sqrt{7}}{7}"


(c)

"\dfrac{x^2}{7}+\dfrac{y^2}{14}=1"

"x=\sqrt{7}\cos t, y=\sqrt{14}\sin t, 0\leq t\leq 2\pi"

"\vec r=\langle\sqrt{7}\cos t, \sqrt{14}\sin t\rangle, 0\leq t\leq 2\pi"

"a=\sqrt{14}, b=\sqrt{7}, c=\sqrt{a^2-b^2}=\sqrt{(\sqrt{14})^2-(\sqrt{7})^2}=\sqrt{7}""e=\dfrac{c}{a}=\dfrac{\sqrt{2}}{2}"

"Foci: (-\sqrt{7}, 0), (\sqrt{7}, 0)"

"Directrices: y=\pm\dfrac{a^2}{c}"

"first\ directrix: y=-2\sqrt{7},"

"second\ directrix:y=2\sqrt{7}"